OpenStudy (anonymous):

It takes Morgan 8 1/3 minutes to run one mile.It takes Kayla 1 1/5 times longer. A.How long does it take Kate to run one mile B. who runs faster C. How long would it take Morgan to run 3 miles

OpenStudy (anonymous):

@Hero

OpenStudy (anonymous):

@myininaya @mathslover

OpenStudy (anonymous):

I would first translate you minutes into seconds what is 8 1/3 in seconds and what is 1 1/5 in seconds

OpenStudy (anonymous):

actually that's not necessary

OpenStudy (anonymous):

@GABI_J do you get the question

OpenStudy (anonymous):

First turn the fractions into decimals

OpenStudy (anonymous):

and you get .3 repeating, and .2

OpenStudy (anonymous):

yes, I thought it would be easier to translate that in seconds first.....

OpenStudy (anonymous):

so 8.3 repeating *1.2

OpenStudy (anonymous):

9.99999999999999 so lets say 10 minutes

OpenStudy (anonymous):

Morgan runs faster

OpenStudy (anonymous):

@Loopki40 tis is workin with multiplying fractions

OpenStudy (anonymous):

Ohh then find the lowest common denominater

OpenStudy (mertsj):

Where you wrote "Kate" did you mean Kayla?

OpenStudy (anonymous):

which is 15, so 1/3 = 5/15 and 1/5 = 3/15, then add them to get 8/15

OpenStudy (anonymous):

Can you help me work out with a

OpenStudy (anonymous):

Yes @Mertsj

OpenStudy (mertsj):

To find Kayla's time: it is 1 1/5 times as long as Morgan so multiply 8 1/3 by 1 1/5

OpenStudy (mertsj):

morgan is faster because her time is less.

OpenStudy (anonymous):

Is that to find a

OpenStudy (mertsj):

Which one asks you to find Kayla's time?

OpenStudy (anonymous):

A

OpenStudy (mertsj):

Then the one that tells you how to find kayla's time is a

OpenStudy (anonymous):

so multiply the 2 numbers to get A

OpenStudy (anonymous):

How do you find b

OpenStudy (mertsj):

If one dress costs $10, how much would 3 dresses cost?

OpenStudy (anonymous):

30

OpenStudy (mertsj):

You should say " 3 times 10" Then you can apply the same thinking: If 1 mile takes 8 1/3 minutes, how long would 3 miles take?

OpenStudy (anonymous):

So I multiply 8 1/3 with 3

OpenStudy (mertsj):

yes

OpenStudy (anonymous):

Can you draw so I can see how to do it

OpenStudy (mertsj):

You mean you don't know how to multiply 8 1/3 by 3?

OpenStudy (anonymous):

Yep to multiply it wouldn't I have to put 3 over 1

OpenStudy (mertsj):

Can you change 8 1/3 to an improper fraction?

OpenStudy (anonymous):

Yes

OpenStudy (mertsj):

So yes. Write 3/1 times 25/3

OpenStudy (mertsj):

Cancel the 3's

OpenStudy (anonymous):

I have a question for a do I change the fractions to improper

OpenStudy (mertsj):

yes

OpenStudy (mertsj):